# Splines-Based Feature Importance in Kolmogorov-Arnold Networks: A Framework for Supervised Tabular Data Dimensionality Reduction

### v1

@@ 4.4 Assessment Procedure (From Feature Selection to Prediction)
4.5 Comparison with Baseline Methods
We compare our spline-based feature selectors against Mutual Information (MI) ranking (Peng et al., 2005), Random Forest (RF) (Breiman, 2001), LASSO-based feature selection (Tibshirani, 1996b), and Recursive Feature Elimination (RFE) for Support Vector Machine (SVM) (SVM-RFE)(Guyon et al., 2002).
5 Experiments
5.1 Data Sets
To evaluate the effectiveness of the Kolmogorov-Arnold Networks dimensinality reduction methods, experiments were conducted on a diverse collection of benchmark datasets covering both classification and regression tasks. These datasets includes well-established repositories from scikit-learn and seaborn as well as synthetically generated datasets and are summarized in Table 1.
Table 1: Benchmark and synthetic datasets
Task Dataset Observations Numbers of features References
Classification Wine 178 13 (Forina et al., 1991).
Breast Cancer Wisconsin 569 30 (Street et al., 1993)
Digits 11,797 64 (Pedregosa et al., 2011)
Make classification 500 10 (Pedregosa et al., 2011)
Regression California Housing 20,640 8 (Efron et al., 2004)
Diabetes 442 10 (Efron et al., 2004).
Diamonds 53,940 10 (Wickham, 2008)
Make regression 500 10 (Pedregosa et al., 2011)
5.2 Results
The comparative analysis across datasets reveals that averaging the macro F1 and $R^{2}$ scores across retention levels $k$ reveals consistent yet dataset-specific interactions between selectors, predictors, and the underlying data structure.
Regression datasets
On the California dataset (nonlinear, heterogeneous), Figure. 1, KAN-L2 and KAN-SI preserve tree-ensemble performance: with Random Forest and XGBoost they stay close to the prediction using all Features (minor drops $\sim$0.01-0.03), while Gradient Boosted Tree is slightly more sensitive. In contrast, KAN-L1 is overly aggressive: it depresses Ridge markedly (large $R^{2}$ gap vs. All Features) and also trims XGBoost and Gradient Boosted Tree more than other KAN variants, indicating loss of weak but complementary signals. KAN-KO feature selector is generally competitive, trailing All Features by a small margin and often matching the classical Mutual Info, Random Forest and SVM-RFE subsets.
Figure 1: Average $R^{2}$ Score (averaged across 20/40/60% retained features) per Selectors for Each Regressor
For Diabetes dataset (small-$n$, noisy), Figure. 1), KAN-L1 again under-selects (lowest bars for Gradient Boosted Tree, Random Forest, XGBoost), whereas KAN-L2, KAN-SI yield the most stable KAN performance across predictors and typically track LASSO and Mutual Info. KAN-KO behaves as a conservative KAN subset, usually second-best among KANs. In the Diamond dataset (strong signal, structured, mixed types), Figure. 1, All models are high-$R^{2}$; here KAN-KO, KAN-L2 and KAN-SI are nearly indistinguishable from All Features for Random Forest, XGBoost and Gradient Boosted Tree and even outperform classical features selectors (Mutual Info, Random Forest and SVM-RFE subsets, LASSO and Random Forest). KAN-L1 is the only KAN variant that noticeably drops on some predictors, consistent with over-pruning informative but correlated attributes. For the Synthetic make regression dataset (well-conditioned, Figure. 1, All KAN variants cluster tightly with baselines across predictors (tiny spreads): when the signal is clean and features are already informative, KAN selection neither helps nor hurts much; KAN-L2/KAN-SI remain the safest choices.
Classification datasets
In the high-dimensional Breast Cancer dataset (Figure 2), Gradient Boosted Trees perform best with KAN-L1, LASSO, and RF importance, often matching the full feature baseline. In contrast, KAN-KO and Mutual Information yield poor subsets, while Logistic Regression and tree ensembles (Random Forest, XGBoost) benefit most from sparsity or impurity-based selectors. This indicates that selectors capturing either linear sparsity or tree-based splits are most effective, whereas knock-out and sensitivity KAN variants are less suited in higher dimensions.
Figure 2: Average Macro $F1$ Score (across 20/40/60% retained features) per Selectors for Each Classifiers
In the high-dimensional Digits dataset (Figure 2), the pattern shifts: Gradient Boosted Trees and XGBoost achieve higher F1 scores with LASSO, SVM-RFE, and several KAN variants (KAN-KO, KAN-SI, KAN-L1), sometimes surpassing the full-feature baseline. Here, dimensionality reduction removes redundancy and sharpens predictive power. By contrast, KAN-L2 and Mutual Information underperform, reflecting their limits in capturing complex feature interactions in high dimension datasets. For the synthetic make classification dataset (Figure 2), all KAN-based selectors (KAN-L1/L2, KAN-SI, KAN-KO) consistently outperform classical baselines, confirming their advantage in structured, high-signal settings where ground-truth feature relevance extends beyond sparsity or univariate dependence. Finally, on the Wine dataset (Figure 2), the advantage shifts back toward classical selectors: RF importance, LASSO, Mutual Information, and SVM-RFE provide the strongest subsets, while most KAN-based selectors lag behind. Only KAN-SI remains competitive, suggesting that when observations are few and feature relationships are well-structured, simpler selectors aligned with linear sparsity or impurity-based measures are more effective.
5.3 Discussion
Across classification and regression, performance hinges on data properties (dimensionality, noise, correlation) and predictor bias. For classification, linear models (Logistic Regression) pair best with sparse and coefficient-driven selectors (LASSO, KAN-L2), whereas ensembles (RF, XGB) benefit from multi-feature selectors (RF importance, KAN-L1/KAN-SI/KAN-KO, SVM-RFE). In high-dimensional sets (Digits), judicious reduction (LASSO, KAN-L1, SVM-RFE) often outperforms the full feature set, revealing harmful redundancy; on synthetic data, all KAN variants recover ground-truth features and rival or beat baselines. For regression, heterogeneous/noisy data (California, Diabetes) favor KAN-L2 and KAN-SI, which track tree ensembles and Ridge with minimal loss, while KAN-L1 is over, aggressive and hurts boosting/linear models. On structured or strong-signal data (Diamond, synthetic), most selectors tie, but KAN-KO/KAN-L2/KAN-SI consistently match or exceed classical baselines.
6 Conclusion
This work presented the first systematic evaluation of Kolmogorov-Arnold Network (KAN) based feature selection across both classification and regression tabular data benchmarks. The results show that KAN selectors, particularly KAN-L2, KAN-SI, and KAN-KO, offer competitive and often superior performance to classical baselines in structured or synthetic settings, while remaining robust on real-world datasets. KAN-L1, though occasionally (for classification datasets) strong, tends to over-prune and degrade predictive power in noisy or correlated contexts especially in regression datasets. Classical selectors such as LASSO and Random Forest remain strong contenders, especially when aligned with linear or ensemble predictors, respectively. Overall, the study highlights that no single selector dominates universally, and that the value of KAN-based feature selection lies in its ability to adaptively capture nonlinear and multivariate dependencies that traditional methods may overlook.

### v2

@@ B.3 Leakage-safe cross-validation
Appendix C KAN-Based Selection Interpretability
In this study, we visualize the internal KAN response functions $f_{j}(x_{j})$ for the three top features on the Breast Cancer dataset Street et al. [1993], worst concave points, mean concave points, and radius error, as identified by KAN-L2, together with the corresponding malignant-class logit. This dataset is based on digitized images of fine-needle aspirates of breast masses, and the prediction task is to distinguish malignant from benign tumours from cell-nucleus morphology. Here, concave points quantify how many segments of a nucleus boundary are inwardly curved (concave), so mean concave points is the average number of such concave segments across all nuclei in an image, and worst concave points is the largest value observed among them. The radius error feature measures the variability of the nucleus radius (standard error of the mean radius) across measurements, capturing irregular or heterogeneous tumour shapes. By plotting the learned responses for these features and the resulting malignant-class logit, we can see how changes in concavity and boundary irregularity are transformed inside the network and how they ultimately influence the predicted malignancy score.
C.1 Mean Concave Points Interpretability Study
For mean concave points, the first-layer KAN response in figure 3(a) is roughly V-shaped. The total contribution (blue curve) is most negative around the centre of the normalized range and rises toward positive values at both low and high extremes. The base term (orange) varies smoothly and almost linearly, while the spline term (green) introduces most of the curvature, indicating that the non-linear spline component is responsible for emphasizing intermediate levels of average concavity in the hidden representation.
When this representation is propagated through the second KAN layer, the malignant-class logit in figure 4(b) becomes strongly skewed. Starting from low mean concave points (very smooth boundaries), the logit is moderately negative (the model leans benign), increases to a clear maximum around the middle of the range (where the model assigns the highest malignancy risk), and then drops sharply for very high values of mean concave points. Thus, in this dataset the KAN regards lesions with intermediate average boundary concavity as most suspicious, while very smooth or extremely concave boundaries are treated as lower risk. Together, the two plots illustrate how a V-shaped internal code for this feature is mapped by the final layer into a peaked malignancy score that is highest at intermediate concavity levels.
(a) KAN first-layer responses of mean concave points
(b) Malignant-class logit as a function of mean concave points (normalized), obtained by varying only this feature around a reference input.
Figure 3: Breast Cancer dataset interpretability visualizations using mean concave points.
C.2 Worst Concave Points Interpretability Study
For worst concave points, the KAN layer response is monotonically decreasing (consistently with the base and spline variation), so increasing concavity systematically pushes the representation along a single direction, in line with the clinical view that strongly concave, spiculated margins are characteristic of malignant lesions. To relate this internal behaviour to the actual prediction, we also examine the malignant-class logit as a function of worst concave points (Figure 4). As worst concavity increases, the malignant logit rises sharply before saturating at extreme values, showing that the internal direction induced by high concavity is ultimately mapped to a higher malignancy score. Together, the first-layer responses and the logit curve indicate that KAN encodes highly concave, irregular tumour margins as a strong, monotone risk factor for malignancy, consistent with established radiological criteria.
(a) KAN first-layer responses of worst concave points
(b) Malignant-class logit as a function of worst concave points (normalized), obtained by varying only this feature around a reference input.
Figure 4: Breast Cancer dataset interpretability visualizations using KAN-L2 feature choice.
C.3 Radius Error Points Interpretability Study
For radius error, the first-layer KAN response is approximately V-shaped (Figure 5(a)), indicating that the network uses a nonlinear code in which moderate values drive the hidden activation most strongly negative, while very small or very large errors have a weaker effect. When this representation is propagated through the second KAN layer, the resulting malignant-class logit becomes roughly monotone in radius error (Figure 5(b)): higher radius error leads to a larger malignant logit with a mild saturation effect. This suggests that the model combines the V-shaped hidden code with other directions in feature space so that, at the prediction level, increasingly irregular tumour borders are treated as a stronger risk factor for malignancy, consistent with radiological practice.
(a) KAN first-layer responses of radius error
(b) Malignant-class logit as a function of radius error points (normalized), obtained by varying only this feature around a reference input.
Figure 5: Breast Cancer dataset interpretability visualizations. (a) KAN first-layer responses for the top three features. (b) Logit sensitivity curve with respect to worst concave points.
Appendix D Stability and Redundancy Analysis
To further assess the reliability of the proposed KAN-L2 feature selection method, we conducted two targeted diagnostic checks on representative datasets.
D.1 Stability
Evaluating the reproducibility of a feature selection method is a central principle in modern feature selection, formalized in stability selection frameworks [Meinshausen and Bühlmann, 2010]. To quantify how consistently each method selects the same features under small perturbations of the dataset, we measure stability across $K=5$ cross-validation folds. For a dataset with feature matrix $X\in\mathbb{R}^{n\times d}$, let $s^{(f)}=(s^{(f)}_{1},\dots,s^{(f)}_{d})$ denote the feature-importance vector computed in fold $f$, and let $\alpha=0.4$ be the retention fraction. We define the selected feature set in fold $f$ as
$A^{(f)}=\operatorname{TopK}\!\left(s^{(f)},\,k\right),\qquad k=\lfloor\alpha d\rfloor.$
To measure agreement between folds $i$ and $j$, we use the Jaccard similarity, a standard stability index in feature selection studies [Nogueira et al., 2018],
$J\!\left(A^{(i)},A^{(j)}\right)=\frac{\lvert A^{(i)}\cap A^{(j)}\rvert}{\lvert A^{(i)}\cup A^{(j)}\rvert},\quad J\in[0,1].$
The overall stability of a method $M$ is given by the mean Jaccard similarity across all $\binom{K}{2}$ fold pairs:
$\mu_{M}=\frac{2}{K(K-1)}\sum_{1\leq i<j\leq K}J\!\left(A^{(i)},\,A^{(j)}\right),$
with variability quantified by the standard deviation
$\sigma_{M}=\sqrt{\frac{2}{K(K-1)}\sum_{1\leq i<j\leq K}\Bigl(J\!\left(A^{(i)},\,A^{(j)}\right)-\mu_{M}\Bigr)^{2}}.$
Higher values of $\mu_{M}$ indicate greater reproducibility of the selected features, while lower $\sigma_{M}$ reflects reduced sensitivity to sampling noise.
Across all four variants (KAN-L1, KAN-KO, KAN-SI, and KAN-L2)(See Figure 6, the stability heatmaps show a consistent pattern: feature selection is perfectly stable on the diamonds dataset (Jaccard $\approx 1.00$for every method), indicating that all selectors repeatedly identify essentially the same subset of features across folds. On the more challenging digits dataset, stability decreases as expected because the feature space is larger and more redundant. Still, the KAN-based methods remain competitive with classical baselines: KAN-L1 ($0.85$), KAN-KO ($0.49$), KAN-SI ($0.72$), and KAN-L2 ($0.62$) all match or exceed the stability of Random Forest permutation importance ($\approx 0.45$), and are generally close to LASSO/L1 ($\approx 0.86$). Overall, these results show that the proposed KAN-based feature scores produce stable and reproducible feature subsets, especially when compared to existing selectors, and remain reliable even in high-dimensional settings
Figure 6: Stability analysis
D.2 Redundancy
To determine whether a method selects mutually correlated (and therefore potentially redundant) predictors, we compare the correlation structure of the full feature space with that of the selected subset. This follows the classical redundancy perspective underlying mRMR-style criteria [Peng et al., 2005]. Let $C=\mathrm{corr}(X)\in\mathbb{R}^{d\times d}$ be the Pearson correlation matrix of all features. The baseline redundancy level is defined as the average absolute correlation across all off-diagonal pairs:
$\overline{\lvert\rho\rvert}_{\text{all}}=\frac{2}{d(d-1)}\sum_{1\leq i<j\leq d}\lvert C_{ij}\rvert$.
For the selected subset $S=A^{(f)}$ of size $k$, the redundancy among selected features is computed from the submatrix $C_{S}=C[S,S]$:
$\overline{\lvert\rho\rvert}_{\text{sel}}=\frac{2}{k(k-1)}\sum_{\begin{subarray}{c}i,j\in S\\ i<j\end{subarray}}\lvert C_{ij}\rvert$.
A method is said to reduce redundancy when
$\overline{\lvert\rho\rvert}_{\text{sel}}<\overline{\lvert\rho\rvert}_{\text{all}},$
indicating that the selected predictors are less correlated than the dataset as a whole. Comparable values suggest that the selector does not inflate redundant structure, even without an explicit sparsity or decorrelation constraint.
Figure 7: Redundancy analysis
KAN variants offer complementary advantages: KAN-L2 provides the cleanest, least-redundant subsets; KAN-L1 yields sparse, efficient selections; KAN-KO balances structure and redundancy; and KAN-SI captures the strongest predictive signals even when features are correlated. Together, they offer flexible, task-adapted feature selection capabilities
Across both datasets in Figure 7, we determined the average feature correlation (all feature vs 40% selected features). The redundancy analysis shows that the KAN feature selectors behave differently depending on the selector. For diamonds, all methods start from the same baseline average correlation ($\approx 0.267$), but their selected subsets diverge significantly: KAN-L2 achieves the strongest redundancy reduction (down to $0.054$), KAN-L1 and KAN-KO reduce redundancy moderately (0.0540.0970.054-0.097), whereas KAN-SI increases redundancy ($0.328$), meaning its top-ranked features are more mutually correlated. For digits, where the baseline redundancy is lower ($\approx 0.124$), KAN-L2 again maintains controlled redundancy ($0.229$), KAN-KO and KAN-SI show modest increases (0.1660.1950.166-0.195), while KAN-SI increases correlation the most. Overall, the results show that KAN-L2 consistently provides the least redundant subsets, KAN-L1 and KAN-KO give moderate redundancy behaviour, and KAN-SI tends to prioritise stronger, more correlated signals, which may help accuracy but does not enforce redundancy minimisation.
Appendix E Prediction Performances per feature retention levels
We provided a predictive analysis in this part for different retention levels.
E.1 Average predictors performances at different retention levels
Figure (8) shows that Across both the Breast Cancer and Digits datasets, the heatmaps show that most selectors achieve near-saturated predictive performance once a moderate fraction of features is retained. For Breast Cancer, accuracy becomes very stable around $94$-$96\%$ from roughly $30$-$40\%$ feature retention onward, indicating substantial redundancy in the original predictors: using only one third of the features is almost as good as using them all. Differences between methods are most visible under aggressive pruning (10-20% retention). In this regime, LASSO/L1, Mutual Information, SVM-RFE, and the KAN-based selectors (especially KAN-L1 and KAN-SI) remain comparatively robust, whereas KAN-KO is noticeably unstable at 10% retention.
Figure 8: Average Accuracy Score (averaged across XGB, Random Forest and Gradient Boosted Trees, and Logistic Regression Classification) for 10/20/30/40/50/60/70/80/90% retained features) for the Digits and Breast cancer data sets
E.2 Insight at 60% retention Level
In this section, we provide a comprehensive overview of the performance results of our models under different feature selection methods across all datasets considered in this study. The appendix is organized into two main subsections: the first subsection focuses on the classification datasets, while the second subsection presents the results for the regression datasets. For each dataset, we report model performance at 60% feature retention, highlighting the impact of feature selection techniques on the predictive accuracy.
E.2.1 Classification data sets
This subsection presents the detailed results for all classification datasets, including breast cancer, digits, make_classification, and wine. Each table reports the macro F1 score of different predictor models using the full features data as well as reduced (60% reduction) selected feature sets obtained via KAN-based selection, LASSO, Mutual Information, Permutation importance, and SVM-RFE. The results illustrate how feature selection affects model performance and help identify the most effective selection strategies for each dataset.
Table 2: 60% retention level for the breast cancer dataset
Models All Features KAN-KO KAN-L1 KAN-L2 KAN-SI LASSO/L1 Mutual Info Perm. (RF) SVM-RFE
GB 0.9624 0.9698 0.9603 0.9603 0.9623 0.9603 0.9623 0.9605 0.9587
LogReg 0.9487 0.9431 0.9433 0.9489 0.9374 0.9488 0.9508 0.9376 0.9526
RF 0.9642 0.9678 0.9698 0.9661 0.9585 0.9678 0.9606 0.9584 0.9547
XGB 0.9716 0.9753 0.9773 0.9754 0.9698 0.9678 0.9718 0.9773 0.9681
Breast cancer. (Table  2) Tree ensembles (GB, RF, XGB) remain competitive or improve with KAN-based selectors. In particular, KAN-L1 attains the top or tied-top accuracy for RF and XGB, and KAN-KO is best for GB. Logistic regression (LogReg) favors sparsity-oriented embedded/wrapper methods (SVM-RFE, LASSO), consistent with its linear inductive bias.
Table 3: 60% retention level for the digits dataset
Models All Features KAN-KO KAN-L1 KAN-L2 KAN-SI LASSO/L1 Mutual Info Perm. (RF) SVM-RFE
GB 0.9869 0.9907 0.9925 0.9813 0.9925 0.9906 0.9907 0.9925 0.9907
LogReg 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 0.9981 1.0000
RF 0.9944 0.9981 1.0000 0.9925 0.9981 0.9962 0.9962 0.9962 0.9963
XGB 0.9963 0.9963 0.9981 0.9926 0.9944 0.9981 0.9981 0.9963 0.9963
Digits. (Table  3) Performance is near a ceiling (LogReg reaches $1.00$), so differences are small. Still, KAN-L1 (and KAN-SI for GB) match or set the best scores for GB/RF/XGB, showing that KAN selectors can remove 40% of inputs without harming multiclass performance on image-like digits.
Table 4: 60% retention level for the make_classification dataset
Models All Features KAN-KO KAN-L1 KAN-L2 KAN-SI LASSO/L1 Mutual Info Permutation (RF) SVM-RFE
GB 0.8838 0.8879 0.8879 0.8879 0.8919 0.8617 0.8819 0.8879 0.8819
LogReg 0.8577 0.8596 0.8596 0.8596 0.8596 0.8477 0.8577 0.8596 0.8577
RF 0.8999 0.9139 0.9059 0.9139 0.9059 0.8819 0.8879 0.9098 0.8919
XGB 0.8959 0.8979 0.9019 0.9039 0.8999 0.8779 0.8959 0.8979 0.8959
Synthetic make_classification. (Table  4) Moderate gains appear for RF (KAN-KO/L2) and GB (KAN-SI), suggesting that removing weak/noisy variables helps tree ensembles; LogReg again prefers sparse selectors (all strong and tied within noise).
Table 5: 60% retention level for the wine dataset
Models All Features KAN-KO KAN-L1 KAN-L2 KAN-SI LASSO/L1 Mutual Info Permutation (RF) SVM-RFE
GB 0.9452 0.9449 0.9566 0.9566 0.9458 0.9505 0.9507 0.9566 0.9450
LogReg 0.9549 0.9663 0.9611 0.9611 0.9610 0.9774 0.9663 0.9611 0.9611
RF 0.9832 0.9837 0.9676 0.9676 0.9784 0.9781 0.9729 0.9728 0.9732
XGB 0.9724 0.9674 0.9615 0.9615 0.9671 0.9667 0.9557 0.9670 0.9728
Wine. (Table  5) Mixed but consistent story: KAN-L1/L2 and Permutation (RF) yield the best or tied-best for GB/XGB, while LogReg peaks with LASSO (as expected). RF is very strong overall and slightly benefits from KAN-KO.
Takeaways. KAN-L1/L2 are reliable selectors for nonlinear learners (GB, RF, XGB), often matching or exceeding the full-feature baseline at 60% retention. KAN-KO is competitive when measured against the end-to-end loss (notably GB and RF), validating a perturbation view of importance. Linear LogReg benefits most from coefficient-based sparsity (LASSO, SVM-RFE). The ability to prune $\approx 40\%$ of features with no loss (and occasional gains) indicates substantial redundancy and supports using KAN-based selection as a practical dimensionality reduction step for tabular classification.
E.2.2 Regression data sets
This subsection presents the results for all regression datasets, including California housing, diabetes, and make_regression. Each table reports the predictive performance of predictor models, measured in terms of $R^{2}$, under various feature selection methods using a data resulting in 60% feature retention, as well as the data having all features. The analysis emphasizes the robustness of models to feature reduction and highlights which selection methods preserve or enhance model performance.
Table 6: 60% retention level for the California dataset
Models All Features KAN-KO KAN-L1 KAN-L2 KAN-SI LASSO/L1 Mutual Info Perm. (RF) SVM-RFE
GB 0.7858 0.7825 0.7416 0.7471 0.7833 0.7565 0.7832 0.7909 0.7563
RF 0.8103 0.8117 0.8266 0.8230 0.8180 0.8257 0.8179 0.8109 0.8257
Ridge 0.6037 0.5886 0.3068 0.3933 0.5854 0.5933 0.5854 0.5946 0.5933
XGB 0.8431 0.8317 0.8318 0.8295 0.8409 0.8284 0.8395 0.8382 0.8261
On California dataset Table 6, ensembles remain resilient at $60\%$ retention. GB peaks with RF permutation; KAN-SI and KAN-KO track the all-features baseline closely. RF improves with KAN-L1, indicating redundancy removal helps tree splits. Ridge is fragile: all-features wins while KAN-L1/L2 over-prune complementary linear signal. XGB prefers all-features, with KAN-SI and Mutual Information nearly matching, SVM-RFE trailing slightly. Across selectors, differences remain modest.
Table 7: 60% retention level for the diabetes dataset
Models All Features KAN-KO KAN-L1 KAN-L2 KAN-SI LASSO/L1 Mutual Info Perm. (RF) SVM-RFE
GB 0.4208 0.4157 0.2405 0.4204 0.4021 0.4326 0.3933 0.4041 0.4205
RF 0.4266 0.4032 0.2601 0.4015 0.4089 0.4132 0.3916 0.4091 0.4186
Ridge 0.4205 0.4209 0.2769 0.4027 0.4127 0.4130 0.4137 0.4127 0.4132
XGB 0.3591 0.3413 0.1716 0.3636 0.3432 0.3554 0.3255 0.3366 0.3682
Diabetes dataset Table 7 exemplifies small-n, noisy regression. GB favors LASSO; RF performs best with all features. Ridge is exceptionally steady near 0.41, tying All and KAN-KO, and barely changing under other selectors. XGB prefers RF permutation. KAN-L1 consistently underperforms, suggesting over-sparsification removes weak yet useful signal. Overall, cautious selection helps, but aggressive pruning hurts most predictors. Mutual Information lags across models slightly.
Diamonds dataset 8 shows a strong signal, near-ceiling performance. Ensembles (RF, GB, XGB) remain virtually unchanged; tiny gains appear with KAN-KO/L2/SI. Ridge benefits modestly from LASSO, suggesting linear redundancy. Overall, selection is largely neutral: most approaches match the all-features baseline within the third decimal. Thus, when structure and signal to noise ratio are high, compact subsets neither help nor harm meaningfully. KAN methods remain competitive
Table 8: 60% retention level for the diamonds dataset
Models All Features KAN-KO KAN-L1 KAN-L2 KAN-SI LASSO/L1 Mutual Info Perm. (RF) SVM-RFE
GB 0.9710 0.9710 0.9708 0.9710 0.9710 0.9704 0.9707 0.9707 0.9707
RF 0.9800 0.9801 0.9792 0.9801 0.9801 0.9800 0.9800 0.9800 0.9800
Ridge 0.8805 0.8806 0.8824 0.8806 0.8806 0.8847 0.8797 0.8797 0.8797
XGB 0.9816 0.9817 0.9810 0.9817 0.9817 0.9810 0.9811 0.9812 0.9811
On synthetic make_regression (Table.  9), signal to noise is high. Ridge sits essentially at one across most selectors, confirming linear recoverability. RF improves modestly with KAN-L2/SI; GB and XGB notch their best with SVM-RFE, though gaps are tiny. KAN variants are stable and competitive, neither overshooting nor collapsing. Overall, many selectors converge, reflecting informative features and limited redundancy. Permutation, LASSO, Mutual Information perform similarly
Table 9: 60% retention level for the make_regression dataset
Models All Features KAN-KO KAN-L1 KAN-L2 KAN-SI LASSO/L1 Mutual Info Permutation (RF) SVM-RFE
GB 0.9456 0.9418 0.9418 0.9501 0.9501 0.9500 0.9499 0.9500 0.9503
RF 0.8819 0.8922 0.8922 0.8976 0.8976 0.8972 0.8972 0.8972 0.8971
Ridge 0.99999 0.9885 0.9885 0.99999 0.99999 0.99999 0.99999 0.99999 0.99999
XGB 0.9233 0.9234 0.9234 0.9325 0.9325 0.9336 0.9341 0.9336 0.9344
Takeaways. KAN-SI and KAN-KO are stable selectors across regression tasks, often matching the all-feature baseline (California, make_regression). KAN-L1/L2 can benefit tree ensembles (RF, GB, XGB) by removing redundancy, but risk over-pruning in noisy, small-$n$ data (Diabetes). Ridge regression is fragile to sparsification, performing best with all features unless signal-to-noise is very high. In high-signal datasets (Diamonds, make_regression), feature selection has negligible effect, with all methods converging near the ceiling. The ability to drop $\approx 40\%$ of features without loss, and sometimes small gains, demonstrates redundancy and supports KAN-based feature selection as a practical dimensionality reduction tool for regression. For the Digits dataset, feature redundancy is even more pronounced. Many selectors already achieve $95$-$98\%$ average accuracy at 10% retention, and by 30-40% retention essentially all methods converge to $\approx 99$-$99.7\%$ accuracy, making them practically indistinguishable. The only clear weakness appears for KAN-L2 at the most extreme pruning level (10%), where performance drops to about $83\%$, but it quickly recovers once more features are kept. Overall, these results suggest that: (i) strong dimensionality reduction is possible without meaningful loss in predictive performance, especially for highly redundant datasets such as Digits; and (ii) KAN-based selectors are competitive with, and often comparable to, classical baselines such as LASSO, Mutual Information, and SVM-RFE, except for a few edge cases under very aggressive pruning.
Appendix F Runtime Profiling of Feature Selectors
Table 10: Selector Runtime Comparison on Diamonds and Wine Datasets
Selector Diamonds (sec) Wine (sec)
Selector-L2 0.022246 0.001013
Selector-L1 0.039464 0.001047
Selector-SI 0.125190 0.003804
LASSO/L1 0.292232 0.001218
Selector-KO 0.426020 0.023628
Mutual Info 2.349357 0.029882
Permutation (RF) 51.709088 0.799606
SVM-RFE 80.940045 0.021224
KAN-Train 527.862040 2.441519
Table 10 reports wall-clock times for computing feature scores on two representative datasets. Once a KAN has been trained, the associated selectors are relatively cheap: coefficient-based criteria (Selector-L1, Selector-L2) are as fast as or faster than LASSO, and the gradient-based score (Selector-SI) remains competitive with mutual information. The knock-out score (Selector-KO) is more expensive, as it requires repeated forward passes with ablated features, but is still considerably lighter than wrapper-style methods such as permutation importance or SVM-RFE, whose runtimes are dominated by repeated model retraining or resampling. The main overhead of our approach lies in the initial KAN training (KAN-Train), which is substantially slower than fitting a single linear or tree model. However, this cost is incurred once per dataset and can then be amortized across multiple selectors, retention levels, and interpretability analyses, whereas wrapper methods must pay a similar price every time a new feature subset is evaluated.